Jack was 14 years older than Priscilla. Together their ages totaled 154 years what were there ages
Answer by stanbon(75887)

Jack was 14 years older than Priscilla. Together their ages totaled 154 years what were there ages
--------
Equations:
j = p + 14
j + p = 154
------
Substitute for "j" and solve for "p":
p+14+p = 154
2p = 140
p = 70 (pricilla'a age)
Slve for "j":
j = p + 14
j = 84 (jack's age)
======================
Cheers,
Stan H.
======================